OpenStudy (anonymous):

Dan had $160 in his bank account. After depositing his paycheck, he had $224 in his bank account. What was the percent of increase in Dan’s bank account after the deposit?

OpenStudy (anonymous):

@Ilovecake

OpenStudy (agl202):

224 -160 = 64 percent = part/whole *100% 64/160 *100 = 640/16 = 40 40% increase

OpenStudy (anonymous):

40% increase

OpenStudy (anonymous):

ok @TheEdwardsFamily

OpenStudy (anonymous):

so

OpenStudy (anonymous):

so it is 40%

OpenStudy (anonymous):

?

OpenStudy (agl202):

Ya!

OpenStudy (anonymous):

@Ilovecake

OpenStudy (anonymous):

yes these are easy anymore?

OpenStudy (ilovecake):

160*(1+x/100)=224

OpenStudy (theedwardsfamily):

40% is right @shadybubba

OpenStudy (ilovecake):

Yes 40%

OpenStudy (anonymous):

ok i got 40%:(

OpenStudy (anonymous):

on my math

OpenStudy (ilovecake):

you got a 40% on a test??

OpenStudy (anonymous):

yes:(

OpenStudy (ilovecake):

I am sooo sorry. Is there anyway I can help?

OpenStudy (anonymous):

na it is all right is is the guy that was helping me t

OpenStudy (theedwardsfamily):

can you tke it again ?

OpenStudy (anonymous):

what is his name

OpenStudy (anonymous):

in 30 minnits

OpenStudy (ilovecake):

I told you that you should round

OpenStudy (anonymous):

yep i shoul have done it

OpenStudy (anonymous):

should

OpenStudy (ilovecake):

I think it was amistre64

OpenStudy (anonymous):

yep

OpenStudy (ilovecake):

that was his name

OpenStudy (anonymous):

i got question 3 and 5 right

OpenStudy (anonymous):

all right im going to eat lunch and in 30 min i will hit you up and we can retake the test ok

OpenStudy (ilovecake):

you got 3 wrong?

OpenStudy (anonymous):

you and TheEdwardsFamily

OpenStudy (ilovecake):

okay. I am sorry :(

OpenStudy (ilovecake):

bye. I am gonna stay online

OpenStudy (theedwardsfamily):

cya

OpenStudy (anonymous):

it is ok we try it one more time and see what happens

OpenStudy (anonymous):

bye

OpenStudy (anonymous):

see you in 30 min

OpenStudy (anonymous):

i will tag you guys in when im reddy

OpenStudy (ilovecake):

k bye